OpenStudy (anonymous):

How do you FOIL x^2-2x+1=9/4

OpenStudy (anonymous):

you have got to be fluttering kidding me man

OpenStudy (anonymous):

It is already foiled...do you mean factor?

OpenStudy (anonymous):

I tried lol I really did. But every time I did, it wouldn't come out the same when I went back to check it.

OpenStudy (anonymous):

motherflutterer feather retricenigger

OpenStudy (anonymous):

Factor x^2-2x+1

OpenStudy (anonymous):

I'm trying to get it where it would look like (x-_)(x-_)=9/4 OR (x+_)(x-_)=9/4 OR however the addition/subtraction sign goes. Now I'm starting to think it doesn't need to be FOIL because of what you said lol :D

OpenStudy (anonymous):

x^2-2x+1 factors to (x-1)^2 or (x-1)(x-1)

OpenStudy (anonymous):

Okay I got it now. You're right, it is (x-1)(x-1). That's what I was looking for. When I did (x-1)(x-1), I didn't get the right answer and I see why now. I made a mistake. Thank you!! :)

OpenStudy (anonymous):

Can you do the rest?

OpenStudy (anonymous):

Let me see if I can, hold on.

OpenStudy (anonymous):

Never mind, I understand it now. I got my final answer. Thank you very much @larryboxaplenty :)